James Corporation is planning to issue bonds with a face value of $504,500 and a coupon rate of 6 percent. The bonds mature in 10 years and pay interest semiannually every June 30 and December 31. All of the bonds will be sold on January 1 of this year. (FV of $1, PV of $1, FVA of $1, and PVA of $1)

Calculate market interest rate (annual) at 4%, 6%, and 8.5%
Answer #1
a) Case B= Interest rate 4%
Bond Price = Present Value Of Interest Payments + Present Value Of Redemption Amount
= P[1-(1+)^-n]/r + Redemption Amount * (1/(1+r)^n
Where,
P= Periodic payment, here interest 504500*6%*6/12=$15135
r= interet rate i.e. 4%*6/12 = 2%
n= number of periods i.e. 10*2= 20
= 15135[1-(1+0.02)^-20]/0.02 + 504500*1/(1+0.02)^20
=247478.94+339514.04
$    5,86,992.98
b) Market Interest Rate 6%
Since bond coupon rate also 6%. The bond issue price would also be equal to face value
=$504500
b) Market Interest Rate 8.5%
Bond Price = Present Value Of Interest Payments + Present Value Of Redemption Amount
= P[1-(1+)^-n]/r + Redemption Amount * (1/(1+r)^n
Where,
P= Periodic payment, here interest 504500*6%*6/12=$15135
r= interet rate i.e. 8.5%*6/12 = 4.25%
n= number of periods i.e. 10*2= 20
= 15135[1-(1+0.0425)^-20]/0.0425 + 504500*1/(1+0.0425)^20
=201210.23+219452.18
$    4,20,662.41